# The Missing Point in Vision Transformers for Universal Image Segmentation

### v1

@@ Abstract
Image segmentation remains a challenging task in computer vision, demanding robust mask generation and precise classification. Recent mask-based approaches yield high-quality masks by capturing global context. However, accurately classifying these masks, especially in the presence of ambiguous boundaries and imbalanced class distributions, remains an open challenge. In this work, we introduce ViT-P, a novel two-stage segmentation framework that decouples mask generation from classification. The first stage employs a proposal generator to produce class-agnostic mask proposals, while the second stage utilizes a point-based classification model built on the Vision Transformer (ViT) to refine predictions by focusing on mask central points. ViT-P serves as a pre-training-free adapter, allowing the integration of various pre-trained vision transformers without modifying their architecture, ensuring adaptability to dense prediction tasks. Furthermore, we demonstrate that coarse and bounding box annotations can effectively enhance classification without requiring additional training on fine annotation datasets, reducing annotation costs while maintaining strong performance. Extensive experiments across COCO, ADE20K, and Cityscapes datasets validate the effectiveness of ViT-P, achieving state-of-the-art results with 54.0 PQ on ADE20K panoptic segmentation, 87.4 mIoU on Cityscapes semantic segmentation, and 63.6 mIoU on ADE20K semantic segmentation. The code and pretrained models are available at https://github.com/sajjad-sh33/ViT-P.
@@ 1 Introduction
Image segmentation involves dividing an image into multiple segments or regions, each representing different objects or parts of objects. This segmentation can be based on semantics, such as distinguishing between buildings, vegetation, and the sky, or on specific instances, such as identifying individual cars, pedestrians, or bicycles within an image. Traditional methods [8, 9, 35, 25] often rely on pixel-wise classification, which struggles to capture the complex relationships between different regions, leading to inaccuracies in object recognition. Recent advancements have led to the development of mask-based segmentation models [15, 14, 27], which significantly improve upon traditional approaches by incorporating global context and spatial dependencies. Instead of classifying each pixel independently, mask-based methods predict a set of binary masks, each corresponding to a single class. This allows for more accurate object separation while preserving spatial coherence.
@@
Mask-based segmentation models, such as Mask2Former [14] and InternImage [46], generate high-quality masks for object instances, enabling more precise segmentation. However, these models often struggle with classifying the generated masks, particularly in scenarios where object boundaries are ambiguous or class distributions are imbalanced. For example, InternImage [46] achieves a mean Intersection-over-Union (mIoU) of 62.9% for semantic segmentation on the ADE20K dataset, but when using ground-truth masks for perfect classification, this improves to 87%. This gap highlights that while state-of-the-art models excel in mask generation, their classification capabilities remain limited across semantic, instance, and panoptic segmentation tasks.
@@
To address these challenges, we introduce a novel two-stage segmentation approach that divides mask generation and classification, enabling more precise segmentation. The first stage employs a proposal generator to produce class-agnostic mask proposals, ensuring robust mask generation. The second stage, ViT-P, utilizes a point-based classification approach, where classification is performed based on the central point of each mask. Built on Vision Transformer (ViT) [18], ViT-P incorporates additional point embeddings into the attention mechanism, allowing simultaneous classification of multiple masks in a computationally efficient manner. ViT-P follows a unified training framework, enabling efficient learning across semantic, instance, and panoptic segmentation tasks without requiring task-specific retraining.
@@
In addition to our decoupled framework, we show that the annotation strategy itself can be optimized for segmentation tasks. While fine-grained annotations are crucial for training robust mask generation modules, they are both time-intensive and expensive to obtain [2]. In contrast, coarse and bounding box annotations can be generated much more efficiently and still provide valuable supervisory signals for mask classification. In our experiments, augmenting the classification stage with coarse and box annotations not only increased classification performance but also allowed us to minimize the overall annotation cost by enabling model training without requiring additional training or pre-training on segmentation datasets. This dual-annotation approach enables us to use a limited number of fine annotations for high-quality mask proposal generation, while taking advantage of abundant, low-cost coarse data to boost classification accuracy.
@@
We evaluated ViT-P across three fundamental image segmentation tasks: panoptic, instance, and semantic segmentation, using three widely adopted benchmarks: COCO [31], Cityscapes [16], and ADE20K [56]. Our proposed architecture demonstrates competitive or superior performance compared to previous state-of-the-art models, highlighting its effectiveness in enhancing mask classification accuracy across diverse segmentation challenges. Specifically, ViT-P improves semantic segmentation results by +0.7% on COCO [31] and 1.3% on ADE20K [56] while utilizing OneFormer [27] for mask generation. Additionally, when paired with InternImage [46], ViT-P boosts Cityscapes [16] semantic segmentation accuracy by +0.4% mIoU. To summarize, our main contributions are:
@@ 2.1 Image Segmentation
In recent years, image segmentation has experienced significant advancements, with various models addressing semantic, instance, and panoptic segmentation challenges. Initially, segmentation was approached as a pixel classification problem, as seen in earlier works such as [8, 9, 12, 25, 35]. However, more recent models like MaskFormer [15] and Mask2Former [14] redefined the task by treating segmentation as a mask classification problem. These models employed a transformer decoder with object queries [5] to generate discrete masks. Progress advanced with models like [26, 45, 4, 11, 20, 46], each building upon the foundations established by Mask2Former [14] and significantly enhancing the field. VLTSeg [26] introduced a novel vision-language pre-training approach, while EVA [20] leveraged masked image modeling, both of which contributed to a more robust segmentation backbone. ViT-Adapter [11] adapted vision transformers to existing frameworks, and InternImage [46] set new benchmarks by employing a novel dynamic sparse convolution framework. ViT-CoMer [47] refined long-range feature interactions, improving mask-based segmentation accuracy. These advancements have set new standards in image segmentation, showcasing the potential of mask-based segmentation approaches for improved accuracy and efficiency.
@@ 2.2 Universal Image Segmentation
Further advancements in image segmentation have led to the development of universal models that unify semantic, instance, and panoptic tasks within a single framework. The aim is to create a model capable of handling diverse tasks without relying on specialized architectures or separate training processes. Recent approaches [22, 27, 50, 51] focus on improving generalization across multiple tasks or datasets rather than designing separate models. The Mask-DINO [30] model, for example, is a unified transformer-based framework for object detection and segmentation, extending DINO [6] by incorporating a mask prediction branch. OneFormer [27] was the first framework to surpass state-of-the-art on all three segmentation tasks with a single universal model, integrating them within a unified framework. FreeSeg [38] employed one-shot training to optimize an all-in-one model for multiple segmentation tasks, seamlessly handling them in a single architecture. OpenSeeD [55] unified visual concepts across segmentation and detection tasks and employed conditioned mask decoding to improve mask generation from bounding boxes. Inspired by recent works such as Mask-DINO [30] and OpenSeeD [55] that exploit box annotations to enhance segmentation, we propose leveraging not only fine annotations but also box and coarse annotations to train segmentation models.
@@ 2.3 Transformers and Pre-Training
Transformers have revolutionized tasks across multiple domains, including natural language processing, computer vision, and speech recognition. Initially introduced for machine translation [43], the vanilla transformer was later adapted for image classification by ViT [18]. DeiT [42] is an efficient version of ViT [18] that significantly reduces the data requirements for training transformers by incorporating a distillation token, enabling the model to learn from a teacher network. BEiT [1] and MAE [24] further advanced ViT [18] for self-supervised learning through masked image modeling, demonstrating the versatility of its architecture. Methods such as MoCov3 [7], DINO [6], DINOv2 [36], iBOT [57], and BEiTv2 [37] have further refined self-supervised learning by leveraging contrastive learning, masked image modeling, and knowledge distillation, enhancing the generalization and robustness of vision transformers across diverse tasks. The advent of CLIP [39] showcased the strength of supervised vision-language pre-training in zero-shot robustness, leading to further developments like EVA-CLIP [41]. To enhance ViT’s flexibility for segmentation while leveraging advanced pre-training, we propose a pre-training-free adapter that incorporates image priors without modifying the original architecture.
@@ 3 Methodology
In this section, we introduce ViT-P, a universal image mask classification model specifically developed based on vision transformers. Using a central point as an input prompt, ViT-P classifies pre-built masks. We first review a meta-architecture for mask-based classification models and describe how classification is the primary limitation of these segmentation methods. Then, we introduce our improved classifier model, which is key to better segmentation in all semantic, instance, and panoptic tasks. Finally, we describe an efficient inference strategy to benefit from both mask- and point-based classifiers.
@@ 3.1 Mask Segmentation Preliminaries
MaskFormer [15] was the initial model that demonstrated that per-pixel classification is not efficient for image segmentation. This model segments an image by predicting binary masks, each with a corresponding category label. After the introduction of MaskFormer [15], all further state-of-the-art models such as [14, 46, 27, 11] followed the same path and tried to enhance the results by improving different parts of the MaskFormer [15] model while keeping the same overall structure.
@@ 3.2 From Mask Creation to Classification
As previously mentioned, all mask-based segmentation models [15, 46, 30, 20, 11] utilize per-segment embeddings to generate masks and their corresponding classes. Our observations reveal that while these models can produce reasonably accurate masks, their classification performance remains suboptimal, significantly reducing overall segmentation performance. As shown in Figure 1, applying perfect classification accuracy for the generated masks (termed the ‘upper bound’) would significantly improve overall segmentation performance. However, the results reveal a substantial gap between this upper bound and the actual performance of existing models, primarily due to weak classification capabilities. To bridge this gap, we propose integrating a separate classification model into the mask classification task, thereby enhancing overall segmentation accuracy.
By exploring various possibilities, we discovered that classifying specific points within generated masks can significantly enhance overall segmentation performance (see Section 4.4 for a detailed setup). In particular, we found that the point with the highest value in a generated mask serves as a critical reference. Accurate classification of this point ensures the correct categorization of the entire mask. Notably, this highest-value point is often located near the center of the mask, maintaining a distance from the borders, which makes classification easier and more reliable. The upper bound results in Figure 1 were also obtained by classifying the generated masks using the label of the highest value point in the ground truth masks. This insight led us to develop a two-stage segmentation model: 1) the first stage employs a proposal generator to produce class-agnostic mask proposals, and 2) then a point-based classification model is employed to classify each mask proposal via classifying its highest value point. The pipeline for the point-based classification model, ViT-P, is described in the following sections.
@@ 3.3 ViT-P Architecture
Figure 2 provides an overview of the ViT-P model. The overall structure of the ViT-P model is based on the vanilla Vision Transformer [18], with additional point embeddings participating in the attention mechanism. While ViT [18] assigns a category to the input image, the ViT-P model takes various point positions as inputs, generating category labels for each input point. This design enables the model to handle mask proposals simultaneously. Similar to ViT [18], patch embeddings are used to divide an image into smaller patches, embedding them as vectors . Here, represents the number of channels, is the number of patches, and is the dimensionality of the embeddings. Additionally, point embeddings are generated using a point encoder, which consists of a linear layer that transforms input normalized points into vectors with the same dimension as the patch embeddings. The embedding sequence is then constructed by concatenating patch embeddings and point embeddings:
@@ 3.4 Training Strategy
As mentioned earlier, for inference, we utilize masks’ central points for better classification results. However, during training, randomly selected points within the image are used to train the ViT-P model. This enables a more robust learning process. Our training methodology leverages three types of annotations (see Figure 3):
@@
Box Annotations: Additionally, we demonstrate that box annotations can be effectively used for pre-training a mask classifier. In this approach, random points within annotation boxes are selected for training, with each point labeled according to its corresponding box. This method introduces a diverse range of annotations, enabling the model to learn effective mask classification despite the lower spatial precision.
In our evaluations, we incorporate a combination of fine, coarse, and box annotations to highlight the effectiveness of our training strategy in addressing high-quality mask misclassification.
@@ 3.5 Integrating Mask and Point Classifiers
During inference, we found it beneficial to merge the classification probabilities of the ViT-P model, denoted as , with the class probabilities of the mask generator model, referred to as . The class probabilities are obtained by removing the "no object" label from the class predictions . Following prior works [21, 49, 53], we employ a geometric ensemble to fuse the classification scores:
where serves as a balance between the predictions of point and mask classifiers. Finally, the "no object" token is concatenated to the fused classification scores, ensuring compatibility with both instance and panoptic segmentation tasks.

### v2

@@ Abstract
Image segmentation remains a challenging task in computer vision, demanding robust mask generation and precise classification. Recent mask-based approaches yield high-quality masks by capturing global context. However, accurately classifying these masks, especially in the presence of ambiguous boundaries and imbalanced class distributions, remains an open challenge. In this work, we introduce ViT-P, a novel two-stage segmentation framework that decouples mask generation from classification. The first stage employs a proposal generator to produce class-agnostic mask proposals, while the second stage utilizes a point-based classification model built on the Vision Transformer (ViT) to refine predictions. ViT-P serves as a pre-training-free adapter, allowing the integration of various pre-trained vision transformers without modifying their architecture, ensuring adaptability to dense prediction tasks. Furthermore, we demonstrate that coarse and bounding box annotations can effectively enhance classification without requiring additional training on fine annotation datasets, reducing annotation costs while maintaining strong performance. Extensive experiments across COCO, ADE20K, and Cityscapes datasets validate the effectiveness of ViT-P, achieving state-of-the-art results with 54.0 PQ on ADE20K panoptic segmentation, 87.4 mIoU on Cityscapes semantic segmentation, and 63.6 mIoU on ADE20K semantic segmentation. The code and pretrained models are available at https://github.com/sajjad-sh33/ViT-P.
@@ 1 Introduction
Image segmentation involves dividing an image into multiple segments or regions, each representing different objects or parts of objects. This segmentation can be based on semantics, such as distinguishing between buildings, vegetation, and the sky, or on specific instances, such as identifying individual cars, pedestrians, or bicycles within an image. Traditional methods [chen2014semantic, chen2017deeplab, long2015fully, huang2019ccnet] often rely on pixel-wise classification, which struggles to capture the complex relationships between different regions, leading to inaccuracies in object recognition. Recent advancements have led to the development of mask-based segmentation models [cheng2021per, cheng2022masked, jain2023oneformer], which significantly improve upon traditional approaches by incorporating global context and spatial dependencies. Instead of classifying each pixel independently, mask-based methods predict a set of binary masks, each corresponding to a single class. This allows for more accurate object separation while preserving spatial coherence.
@@
Mask-based segmentation models, such as Mask2Former [cheng2022masked] and InternImage [wang2023internimage], generate high-quality masks for object instances, enabling more precise segmentation. However, these models often struggle with classifying the generated masks, particularly in scenarios where object boundaries are ambiguous or class distributions are imbalanced. For example, InternImage [wang2023internimage] achieves a mean Intersection-over-Union (mIoU) of 62.9% for semantic segmentation on the ADE20K dataset, but when using ground-truth masks for perfect classification, this improves to 87%. This gap highlights that while state-of-the-art models excel in mask generation, their classification capabilities remain limited across semantic, instance, and panoptic segmentation tasks.
@@
To address these challenges, we introduce a novel two-stage segmentation approach that divides mask generation and classification, enabling more precise segmentation. The first stage employs a proposal generator to produce class-agnostic mask proposals, ensuring robust mask generation. The second stage, ViT-P, utilizes a point-based classification approach, where classification is performed based on the highest value point of each mask. Built on Vision Transformer (ViT) [dosovitskiy2020image], ViT-P incorporates additional point embeddings into the attention mechanism, allowing simultaneous classification of multiple masks in a computationally efficient manner. ViT-P follows a unified training framework, enabling efficient learning across semantic, instance, and panoptic segmentation tasks without requiring task-specific retraining.
@@
In addition to our decoupled framework, we show that the annotation strategy itself can be optimized for segmentation tasks. While fine-grained annotations are crucial for training robust mask generation modules, they are both time-intensive and expensive to obtain [bearman2016s]. In contrast, coarse and bounding box annotations can be generated much more efficiently and still provide valuable supervisory signals for mask classification. In our experiments, augmenting the classification stage with coarse and box annotations not only increased classification performance but also allowed us to minimize the overall annotation cost by enabling model training without requiring additional training or pre-training on segmentation datasets. This dual-annotation approach enables us to use a limited number of fine annotations for high-quality mask proposal generation, while taking advantage of abundant, low-cost coarse data to boost classification accuracy.
@@
We evaluated ViT-P across three fundamental image segmentation tasks: panoptic, instance, and semantic segmentation, using three widely adopted benchmarks: COCO [lin2014microsoft], Cityscapes [cordts2016cityscapes], and ADE20K [zhou2017scene]. Our proposed architecture demonstrates competitive or superior performance compared to previous state-of-the-art models, highlighting its effectiveness in enhancing mask classification accuracy across diverse segmentation challenges. Specifically, ViT-P improves semantic segmentation results by +0.7% on COCO [lin2014microsoft] and 1.3% on ADE20K [zhou2017scene] while utilizing OneFormer [jain2023oneformer] for mask generation. Additionally, when paired with InternImage [wang2023internimage], ViT-P boosts Cityscapes [cordts2016cityscapes] semantic segmentation accuracy by +0.4% mIoU. To summarize, our main contributions are:
@@
State-of-the-Art Performance Across Benchmarks. ViT-P achieves 54.0 PQ on ADE20K [zhou2017scene] panoptic segmentation, 49.0 AP on Cityscapes [cordts2016cityscapes] instance segmentation, and 53.7 mIoU on COCO-Stuff-164K [caesar2018coco] semantic segmentation, setting new performance benchmarks in segmentation.
@@ 2.1 Image Segmentation
In recent years, image segmentation has experienced significant advancements, with various models addressing semantic, instance, and panoptic segmentation challenges. Initially, segmentation was approached as a pixel classification problem, as seen in earlier works such as [chen2014semantic, chen2017deeplab, cheng2019spgnet, huang2019ccnet, long2015fully]. However, more recent models like MaskFormer [cheng2021per] and Mask2Former [cheng2022masked] redefined the task by treating segmentation as a mask classification problem. These models employed a transformer decoder with object queries [carion2020end] to generate discrete masks. Progress advanced with models like [hummer2024strong, wang2023image, cai2022reversible, chen2022vision, fang2023eva, wang2023internimage], each building upon the foundations established by Mask2Former [cheng2022masked] and significantly enhancing the field. VLTSeg [hummer2024strong] introduced a novel vision-language pre-training approach, while EVA [fang2023eva] leveraged masked image modeling, both of which contributed to a more robust segmentation backbone. ViT-Adapter [chen2022vision] adapted vision transformers to existing frameworks, and InternImage [wang2023internimage] set new benchmarks by employing a novel dynamic sparse convolution framework. ViT-CoMer [xia2403vit] refined long-range feature interactions, improving mask-based segmentation accuracy. These advancements have set new standards in image segmentation, showcasing the potential of mask-based segmentation approaches for improved accuracy and efficiency.
@@ 2.2 Universal Image Segmentation
Further advancements in image segmentation have led to the development of universal models that unify semantic, instance, and panoptic tasks within a single framework. The aim is to create a model capable of handling diverse tasks without relying on specialized architectures or separate training processes. Recent approaches [gu2023dataseg, jain2023oneformer, xu2024rap, yan2023universal] focus on improving generalization across multiple tasks or datasets rather than designing separate models. The Mask-DINO [li2023mask] model, for example, is a unified transformer-based framework for object detection and segmentation, extending DINO [caron2021emerging] by incorporating a mask prediction branch. OneFormer [jain2023oneformer] was the first framework to surpass state-of-the-art on all three segmentation tasks with a single universal model, integrating them within a unified framework. FreeSeg [qin2023freeseg] employed one-shot training to optimize an all-in-one model for multiple segmentation tasks, seamlessly handling them in a single architecture. OpenSeeD [zhang2023simple] unified visual concepts across segmentation and detection tasks and employed conditioned mask decoding to improve mask generation from bounding boxes. Inspired by recent works such as Mask-DINO [li2023mask] and OpenSeeD [zhang2023simple] that exploit box annotations to enhance segmentation, we propose leveraging not only fine annotations but also box and coarse annotations to train segmentation models.
@@ 2.3 Transformers and Pre-Training
Transformers have revolutionized tasks across multiple domains, including natural language processing, computer vision, and speech recognition. Initially introduced for machine translation [vaswani2017attention], the vanilla transformer was later adapted for image classification by ViT [dosovitskiy2020image]. DeiT [touvron2021training] is an efficient version of ViT [dosovitskiy2020image] that significantly reduces the data requirements for training transformers by incorporating a distillation token, enabling the model to learn from a teacher network. BEiT [bao2021beit] and MAE [he2022masked] further advanced ViT [dosovitskiy2020image] for self-supervised learning through masked image modeling, demonstrating the versatility of its architecture. Methods such as MoCov3 [chen2024self], DINO [caron2021emerging], DINOv2 [oquab2023dinov2], iBOT [zhou2021ibot], and BEiTv2 [peng2022beit] have further refined self-supervised learning by leveraging contrastive learning, masked image modeling, and knowledge distillation, enhancing the generalization and robustness of vision transformers across diverse tasks. The advent of CLIP [radford2021learning] showcased the strength of supervised vision-language pre-training in zero-shot robustness, leading to further developments like EVA-CLIP [sun2023eva]. To enhance ViT’s flexibility for segmentation while leveraging advanced pre-training, we propose a pre-training-free adapter that incorporates image priors without modifying the original architecture.
@@ 3 Methodology
In this section, we introduce ViT-P, a universal image mask classification model specifically developed based on vision transformers. Using point positions as an input prompt, ViT-P classifies pre-built masks. We first review a meta-architecture for mask-based classification models and describe how classification is the primary limitation of these segmentation methods. Then, we introduce our improved classifier model, which is key to better segmentation in all semantic, instance, and panoptic tasks. Finally, we describe an efficient inference strategy to benefit from both mask- and point-based classifiers.
@@ 3.1 Mask Segmentation Preliminaries
MaskFormer [cheng2021per] was the initial model that demonstrated that per-pixel classification is not efficient for image segmentation. This model segments an image by predicting binary masks, each with a corresponding category label. After the introduction of MaskFormer [cheng2021per], all further state-of-the-art models such as [cheng2022masked, wang2023internimage, jain2023oneformer, chen2022vision] followed the same path and tried to enhance the results by improving different parts of the MaskFormer [cheng2021per] model while keeping the same overall structure.
@@ 3.2 From Mask Creation to Classification
As previously mentioned, all mask-based segmentation models [cheng2021per, wang2023internimage, li2023mask, fang2023eva, chen2022vision] utilize per-segment embeddings to generate masks and their corresponding classes. Our observations reveal that while these models can produce reasonably accurate masks, their classification performance remains suboptimal, significantly reducing overall segmentation performance. As shown in Figure LABEL:Figure:figure1, applying perfect classification accuracy for the generated masks (termed the ‘upper bound’) would significantly improve overall segmentation performance. However, the results reveal a substantial gap between this upper bound and the actual performance of existing models, primarily due to weak classification capabilities. To bridge this gap, we propose integrating a separate classification model into the mask classification task, thereby enhancing overall segmentation accuracy.
By exploring various possibilities, we discovered that classifying specific points within generated masks can significantly enhance overall segmentation performance (see Section 4.4 for a detailed setup). In particular, we found that the point with the highest value in a generated mask serves as a critical reference. Accurate classification of this point ensures the correct categorization of the entire mask. Notably, this highest-value point is often located near the center of the mask, maintaining a distance from the borders, which makes classification easier and more reliable. The upper bound results in Figure LABEL:Figure:figure1 were also obtained by classifying the generated masks using the label of the highest value point in the ground truth masks. This insight led us to develop a two-stage segmentation model: 1) the first stage employs a proposal generator to produce class-agnostic mask proposals, and 2) then a point-based classification model is employed to classify each mask proposal via classifying its highest value point. The pipeline for the point-based classification model, ViT-P, is described in the following sections.
@@ 3.3 ViT-P Architecture
Figure 2 provides an overview of the ViT-P model. The overall structure of the ViT-P model is based on the vanilla Vision Transformer [dosovitskiy2020image], with additional point embeddings participating in the attention mechanism. While ViT [dosovitskiy2020image] assigns a category to the input image, the ViT-P model takes various point positions as inputs, generating category labels for each input point. This design enables the model to handle mask proposals simultaneously. Similar to ViT [dosovitskiy2020image], patch embeddings are used to divide an image into smaller patches, embedding them as vectors . Here, represents the number of channels, is the number of patches, and is the dimensionality of the embeddings. Additionally, point embeddings are generated using a point encoder, which consists of a linear layer that transforms input normalized points into vectors with the same dimension as the patch embeddings. The embedding sequence is then constructed by concatenating patch embeddings and point embeddings:
@@ 3.4 Training Strategy
As mentioned earlier, for inference, we utilize masks’ highest value points for better classification results. However, during training, randomly selected points within the image are used to train the ViT-P model. This enables a more robust learning process. Our training methodology leverages three types of annotations (see Figure 3):
@@
Box Annotations: Bounding box annotations are utilized exclusively for a pre-training stage. During this phase, we do not sample points. Instead, the model is trained to classify objects using the box coordinates directly as input. A box is represented by its coordinates (e.g., center point, width, and height), which are fed into the ViT-P model.
Our training strategy involves first pre-training the ViT-P model using the efficient box annotations. This teaches the model a general understanding of object location and scale. Subsequently, the model is fine-tuned using the more precise point-based inputs from either fine or coarse annotations. To bridge the gap between the pre-training (box inputs) and fine-tuning (point inputs) stages, we adapt the input format. When fine-tuning, a sampled point with coordinates is passed to the model as . This effectively represents the point as a box with zero height and width, maintaining a consistent input structure for the model. In our evaluations, we demonstrate the effectiveness of this combined strategy, pre-training on boxes and fine-tuning on points, in addressing high-quality mask misclassification.
@@ 3.5 Integrating Mask and Point Classifiers
During inference, we found it beneficial to merge the classification probabilities of the ViT-P model, denoted as , with the class probabilities of the mask generator model, referred to as . The class probabilities are obtained by removing the ”no object” label from the class predictions . Following prior works [ghiasi2022scaling, xu2023open, yu2023convolutions], we employ a geometric ensemble to fuse the classification scores:
where serves as a balance between the predictions of point and mask classifiers. Finally, the ”no object” token is concatenated to the fused classification scores, ensuring compatibility with both instance and panoptic segmentation tasks.
Method #Params #FLOPs PQ AP UPerNet [xiao2018unified] SwinV2-L [liu2022swin] — — 640640 — — — 55.9 UPerNet [xiao2018unified] InternImage-L [wang2023internimage] 256 2526 640640 — — 53.9 54.1 UPerNet [xiao2018unified] InternImage-XL [wang2023internimage] 368 3142 640640 — — 55.0 55.3 MaskFormer [cheng2021per] Swin-L [liu2021swin] 212 375 640640 — — 54.1 55.6 Mask2Former-Panoptic [cheng2022masked] Swin-L [liu2021swin] 216 413 640640 48.7 34.2 54.5 — OneFormer [jain2023oneformer] ConvNeXt-L [liu2022convnet] 220 389 640640 50.0 36.2 56.6 57.4 OneFormer [jain2023oneformer] DiNAT-L [hassani2022dilated] 223 1768 12801280 51.5 37.1 58.3 58.7 OneFormer† [jain2023oneformer] DiNAT-L [hassani2022dilated] 223 1768 12801280 53.4 40.2 58.4 58.8 OneFormer [jain2023oneformer] ConvNeXt-XL [liu2022convnet] 372 607 640640 50.1 36.3 57.4 58.8 OpenSeeD†‡ [zhang2023simple] Swin-L [liu2021swin] 286 — 12801280 53.7 42.6 58.4 — OneFormer+ViT-P (ours) DiNAT-L [hassani2022dilated] 309 1955 12801280 51.9 37.8 58.6 59.0 OneFormer†+ViT-P (ours) DiNAT-L [hassani2022dilated] 309 1955 12801280 54.0 40.7 59.7 59.9 UPerNet [xiao2018unified] SwinV2-G [liu2022swin] 3 — 640640 — — 59.1 — OneFormer [jain2023oneformer] InternImage-H [wang2023internimage] 1.1 4193 896896 54.5 40.2 60.4 60.8 Mask2Former† [cheng2022masked] RevCol-H [cai2022reversible] 2.4 — 640640 — — 60.4 61.0 Mask2Former† [cheng2022masked]+ViT-Adapter [chen2022vision] BEiT-3 [wang2023image] 1.9 — 896896 — — 62.0 62.8 Mask2Former† [cheng2022masked] InternImage-H [wang2023internimage] 1.31 4635 896896 — — 62.5 62.9 OneFormer+ViT-P (ours) InternImage-H [wang2023internimage] 1.4 4812 896896 54.5 40.6 61.2 61.6 Mask2Former†+ViT-P† (ours) InternImage-H [wang2023internimage] 1.61 5254 896896 — — 63.1 63.6
Method Coarse #Params #Queries PQ AP Axial-DeepLab-XL [wang2020axial] Axial ResNet-XL [wang2020axial] 173 — 10252049 64.4 36.7 80.6 81.1 Panoptic-DeepLab [cheng2020panoptic] SWideRNet [chen2020scaling] 536 — 10252049 66.4 40.1 82.2 82.9 Mask2Former-Panoptic [cheng2022masked] Swin-L [liu2021swin] 216 200 5121024 66.6 43.6 82.9 — Mask2Former-Semantic [cheng2022masked] Swin-L [liu2021swin] 215 100 5121024 — — 83.3 84.3 OneFormer [jain2023oneformer] ConvNeXt-L [liu2022convnet] 220 250 5121024 68.5 46.5 83.0 84.0 OneFormer† [jain2023oneformer] ConvNeXt-L [liu2022convnet] 220 250 5121024 70.0 48.7 84.6 85.2 OneFormer† [jain2023oneformer] ConvNeXt-XL [liu2022convnet] 372 250 5121024 69.7 48.9 84.5 85.8 OpenSeeD‡ [zhang2023simple] Swin-L [liu2021swin] 286 300 5121024 68.9 48.5 84.5 — OneFormer†+ViT-P ConvNeXt-L [liu2022convnet] ✓ 306 250 5121024 70.0 49.0 84.9 85.5 OneFormer [jain2023oneformer] InternImage-H [wang2023internimage] 1.1 250 5121024 70.6 50.6 85.1 85.7 Mask2Former [cheng2022masked]+Depth Anything [yang2024depth] ViT-L [dosovitskiy2020image] — 200 896896 — — 84.8 86.2 Mask2Former† [cheng2022masked] InternImage-H [wang2023internimage] 1.1 100 10241024 — — 86.4 87.0 OneFormer+ViT-P InternImage-H [wang2023internimage] ✓ 1.4 250 5121024 70.8 50.6 85.4 85.9 Mask2Former†+ViT-P† InternImage-H [wang2023internimage] ✓ 1.4 100 10241024 — — 86.8 87.4
COCO val2017 COCO-Stuff-164K Method #Params PQ AP MaskFormer [cheng2021per] Swin-L [liu2021swin] 212 640640 52.7 — 64.8 — — Mask2Former-Panoptic [cheng2022masked] Swin-L [liu2021swin] 216 10241024 57.8 48.6 67.4 — — kMaX-DeepLab [yu2022k] ConvNeXt-L [liu2022convnet] 232 12811281 58.0 — — — — OneFormer [jain2023oneformer] Swin-L [liu2021swin] 219 10241024 57.9 49.0 67.4 — — OneFormer [jain2023oneformer] DiNAT-L [hassani2022dilated] 223 10241024 58.0 49.2 68.1 — — OpenSeeD† [zhang2023simple] Swin-L [liu2021swin] 286 10241024 59.5 53.2 68.6 — — OneFormer+ViT-P (ours) DiNAT-L [hassani2022dilated] 309 10241024 58.0 49.5 68.6 — — OneFormer+ViT-P† (ours) DiNAT-L [hassani2022dilated] 309 10241024 58.2 49.6 68.8 — — OneFormer [jain2023oneformer] InternImage-H [wang2023internimage] 1.35 10241024 60.0 52.0 68.8 — — Mask2Former [cheng2022masked]+ViT-Adapter [chen2022vision] BEiTv2 [peng2022beit] 571 896896 — — — 52.3 — Mask2Former [cheng2022masked]+ViT-Adapter [chen2022vision] EVA [fang2023eva] — 896896 — — — 53.4 — Mask2Former [cheng2022masked] InternImage-H [wang2023internimage] 1.31 896896 — — — 52.6 52.8 Mask2Former+ViT-P (ours) InternImage-H [wang2023internimage] 1.61 896896 — — — 53.5 53.7